If E = [ 1/3 + 1/50 ] + [ 1/3 + 2/50 ]+....+ upto 50 terms then what is the exponent of 2 in E ! ?

Grade:10

1 Answers

mycroft holmes
8 years ago
Use Hermite’s identity:
 
Then E =[50/3] = 16.
 
Then the exponent of 2 in 16! = [16/2]+[16/4]+[16/8]+[16/16] = 8+4+2+1 = 15